Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name Annand
Log entry time 16:05:42 on June 8,2009
Entry number 279915
keyword=Day shift summary
Day shift summary: J.R.M.Annand (SL), R.De Leo (TO), F.Garibaldi (3rd)
=============================================================================
Events
------
Continue production running in transverse-minus He3
No projected changfes in configuration
a.m. beam has been rather unstable
Lost beam approx 10.40-11.00 RF cavity tune
Unstable beam conditions persist throughout the shift due to RF
cavity problems
RUNS-LHRS/BB
---------
3351 Period 08.06-08.44 pol.He3, production 4.00M events
3352 Period 08.49-09.23 pol.He3, production 4.00M
3353 Period 09.26-09.56 pol.He3, production 4.00M
3354 Period 09.58-10.29 pol.He3, production 4.00M
3355 Period 10.31-11.00 pol.He3, production 1.62M lost beam during run
3356 Period 11.06-11.36 pol.He3, production 4.00M
3357 Period 11.38-12.11 pol.He3, production 4.00M
3358 Period 12.16-12.50 pol.He3, production 4.00M
3359 Period 12.57-13.30 pol.He3, production 4.00M
3360 Period 13.32-14.27 pol.He3, production 4.00M 20min lost beam during run
3361 Period 14.34-15.07 pol.He3, production 4.00M
3362 Period 15.09-15.54 pol.He3, production 4.00M
3363 Period 15.56 pol.He3 continued in swing
RUNS-RHRS/ND
---------
22224 Period 08.04-08.45 pol.He3, production 2.86M events
22225 Period 08.48-09.24 pol.He3, production 2.85M
22226 Period 09.25-09.56 pol.He3, production 2.85M
22227 Period 09.57-10.29 pol.He3, production 2.86M
22228 Period 10.30-11.00 pol.He3, production 1.16M lost beam during run
22229 Period 11.06-11.36 pol.He3, production 2.85M
22230 Period 11.38-12.12 pol.He3, production 2.87M
22231 Period 12.14-12.51 pol.He3, production 2.85M
22232 Period 12.56-13.30 pol.He3, production 2.86M
22233 Period 13.31-14.27 pol.He3, production 2.88M 20min lost beam during run
22234 Period 14.33-15.07 pol.He3, production 2.85M
22235 Period 15.09-15.54 pol.He3, production 2.87M
22236 Period 15.55 pol.He3, continued in swing
RUNS-Compton
------------
17253 Period 06.30-08.30 end run failed, reset and downloaded on coda gui
17254 Period 08.37-10.30 end run failed, reset and downloaded on coda gui
17255 Period 10.34-12.34 end run failed, reset and downloaded on coda gui
17256 Junk rebooted rocs and restarted coda
17257 Period 12.41-?
Disappeared runs 17258-17260...what's going on?
17261 Period 14.47- rebooted rocs and restarted coda
RUNS-HAPPEX
-----------
31496 Period 04.30-08.45 carried from owl shift
31497 Period 08.45-12.51
31498 Period 12.51- continued into swing shift